How would you solve the following problems:
3(-2-3x)=-9x-4
5/2t-t=3/2t+3

Answer by WannabeCAgirl83(35)

3(-2-3x) = -9x-4 → get rid of the ( ) by multiplying 3 with -2 and -3x
-6-9x = -9x-4 | +6 on both sides
-9x = -9x+2 → You can't bring the x term on either side of the equation because if you added +9x the term would disapper on both sides as -9x +9x = 0
This is unsolvable, I think.

5/2t-t = 3/2t+3 | ∙ 2t on both sides
5-2tē = 3+3 | -5 on both sides
-2tē = 1 | : -2
tē = -0.5 | square root
t1 = 0.707106781 ^ t2 = -0.707106781